OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = –31 and a9 = –119?

OpenStudy (amistre64):

a9 = a1 + d(8) solve for d with the given information to determine the 32nd term

OpenStudy (amistre64):

a32 = a1+d(31)

OpenStudy (anonymous):

whats the anwser?

OpenStudy (amistre64):

what answer id you get when you did that?

OpenStudy (anonymous):

i got -339

OpenStudy (amistre64):

let me chk if thats good :)

OpenStudy (amistre64):

-119 =-31 + d(8) (-119+31)/8 = d a32 = -31 + ((-119+31)/8)(31) = -372

OpenStudy (anonymous):

oh i see thank you !!

OpenStudy (amistre64):

your welcome :)

OpenStudy (anonymous):

is it -339

OpenStudy (anonymous):

No I believe it is -372 like amistre had said...not sure though